Use a graphing utility, where helpful, to find the area of the region enclosed by the curves.
y=x^3−13x^2+30x and y=0.
Carry out all calculations exactly and round to 2 decimal places the final answer only.

Answer by Alan3354(69443)

Use a graphing utility, where helpful, to find the area of the region enclosed by the curves.
y=x^3-13x^2+30x and y=0.
===============
y=x^3-13x^2+30x
INT = x^4/4 - 13x^3/3 + 15x^2 + C
-----
y = 0 is the x-axis
-----
There are 2 zeroes.
x = 3 and x = 10
The area from x = 0 to x = 3 is above the x-axis.
x^4/4 - 13x^3/3 + 15x^2 + C
f(0) = 0
f(3) = 81/4 - 117 + 135
f(3) - f(0) = 153/4 sq units
==============
The area from x=3 to x=10 is below the x-axis.
f(10) = 2500 - 13000/3 + 1500 = -1000/3
f(10) - f(3) = -4459/12 sq units
============================
IDK how you want to combine the two areas.